# Situationally-Aware Dynamics Learning

### v2

@@ V-F2 Real World Experiments:
We evaluate the zero-shot performance of our situationally-aware dynamics model by applying it directly from simulation to real-world scenarios without additional fine-tuning. Figure 14 highlights various challenging terrains successfully navigated using our method, demonstrating its robustness and adaptability in real-world conditions. The accompanying video (https://youtu.be/VKR18WaSCAk) showcases the robot successfully navigating these scenarios.
@@
We also move to more challenging tasks with steeper and more rugged terrain, which is shown in Figure 4 (right). To protect the hardware, we only evaluated the best performing baseline in the previous task (Dubins). However, we had to stop the experiment with the Dubins model as it repeatedly led to the robot turning over and rolling down the mountain, as shown in Figure 16. In contrast, our method repeatedly completed the task, as can be seen starting at the 1:42 minute mark. Overall, our method can lead the robot to successfully climb bumpy and muddy mountains, whereas the baselines fail as they get stuck on terrain irregularities or are unable to devise a safe and successful plan.
@@
  • Backing Up: Whenever forward motion is impeded, we see the robot back up and try moving with a slight orientation change to overcome the terrain irregularities. This mirrors the behavior of human drivers when their vehicles become stuck in challenging terrains. Check the clips at the: 1:00, 1:27 (trial 4), 2:06, 2:36 (trial 4), 3:30 & 5:25 (rocks) minute marks for evidence of this.
  • Velocity Control: We have seen multiple evidence of how the robot’s planned velocity commands differ according to its current situation. For example, in Figure 12 we saw, for every situation, different plans to reach the same goal from the same starting state. Moreover, at the 4:00 minute mark of the accompanying video, the robot maintains a slower pace while navigating downhill and then deliberately accelerates (minute 4:15) upon a situation change–reaching safer, more stable terrain.
  • Trajectory Smoothness: On very rugged and dangerous terrains the robot followed curved patterns (see the clips at minutes 1:42, 2:22 & 3:18). On the other hand, when the terrain was not as rugged and steep, the robot approached its targets directly, relying only on the velocity control to maintain its integrity (4:00, 4:22 & 5:22 minute marks).
@@
We dive deeper into why we believe this is happening in Section VI. At a high-level, we hypothesize that upon a situation change, the optimization process carried out by the sampling-based planner will be bounded by new constraints given by the dynamics of the current situation, thus generating new diverse action plans.
@@
Figure 15: Results from Real World Experiments. We repeat the experiments ten times with each baseline and report the success rate, number of waypoints reached, the cumulative time between successive waypoints and the total time to comple the task.
@@
Figure 16: We test our situationally-aware dynamics model against the physics-based Dubins model (best performing baseline) in the challenging task of descending a steep mountain with wet leaves and many large bumps. Across multiple trials, our model consistently succeeded in completing the task, whereas the baseline failed, causing the robot to flip.
VI Discussion
@@
This study presented a framework for learning online and unsupervised hidden state representations to inform the dynamics models of robots operating in complex, unstructured environments. By integrating these learned representations with the dynamics models, we demonstrated significant performance improvements over traditional approaches that rely solely on raw sensory inputs. Particularly interesting was the emergence of safe behaviors that unlocked a new range of tasks the robot could safely accomplish. Here, we dive into the implications of these results, their connection to prior work, and the broader insights they provide for learning-based dynamics modeling in robotics.
@@
Generalization to Novel Terrains. One of the most encouraging findings was the model’s ability to generalize to unseen terrains directly without additional work from simulation to reality and the emergence of safe navigation behaviors. This suggests that the learned hidden states encode features that are transferable across a broad range of terrain types. Such generalization capabilities are crucial for autonomous systems that must operate in diverse environments without extensive retraining, particularly in scenarios such as disaster response or planetary exploration.
Data Efficiency and Model Scalability. Despite the complexity of the augmented dynamics model, it demonstrated significant data efficiency compared to baseline methods. By leveraging the representations learned online, the model required fewer training episodes to achieve superior performance, as evidenced in Figures 6 & 9. This efficiency highlights the potential for scaling the approach to larger, more complex systems, and even online dynamics learning tasks, since we abandon the need for privileged information and extensive offline training from previous works on auxiliary representation learning.
@@
Integration with Planning and Control. The improved dynamics predictions directly impacted downstream planning and control tasks, leading to safer and adaptive navigation. As shown in Figures 14, 15 & 16, the performance improvement is notably higher in challenging real world scenarios. By augmenting the state space with the hidden representations, the controller could generate trajectories that proactively accounted for terrain-induced dynamics changes.
@@
Comparison with Learning-Based Models. The performance of the CaDM and PE+LSTM baselines highlights the inherent difficulty of learning hidden state representations alongside dynamics in a data-efficient manner. While these methods have shown promise in other scenarios, our online hidden state representation framework, grounded in multivariate BOCD, offers a more effective alternative. By leveraging structured latent spaces and online learning, our approach eliminates the need for interdependent, separately trained models, addressing the problem from a novel perspective. On the dynamics learning side, the PE-Y model, despite using and predicting the same variables as the Dubins model, exhibited poor real-world performance. Similarly, PE-RPY struggled to generalize to new terrains; its requirement to predict the robot’s orientation (roll, pitch, and yaw) introduced compounding errors during model-based planning, which led to suboptimal strategies, as demonstrated in minute 0:51 (PE-RPY) of the video. We further hypothesize that TD-MPC2’s slower learning is attributable to its large model size, necessitating the simultaneous learning of five components (encoder, latent dynamics, reward predictor, terminal value predictor, and policy), which may introduce unnecessary overhead for this specific task. Finally, while the model-free alternatives predictably underperformed in terms of learning efficiency, they underscore the advantages of a model-based approach for adaptive robot behavior. This is particularly compelling when contrasted with the policy-learning strategies for adaptive robot behaviors employed by other works, as discussed in Section II-C.
@@
Comparison with Physics-Based Models. The Dubins model was the most effective baseline but with critical flaws in complex real world scenarios. Note that its performance stems from being designed by an expert with extensive domain knowledge. This adds some limitations due to the stringent prerequisites to replicate this approach for other robots. Additionally, it operates under the assumption of ideal conditions, which limits its applicability in real-world, unstructured terrain navigation tasks. Critically, the model lacks a feedback mechanism to account for the specific characteristics of varying terrains, making it ill-suited for dynamic and complex environments. This limitation was evident in several scenarios where the robot’s performance was unsatisfactory. For instance, the model was unable to compensate for latent factors affecting the robot’s motion—as captured in minute 0:37 (Dubins)—and frequently resulted in unsafe behaviors, both for the robot and its surroundings, as demonstrated in Figure 16.
@@
Broader Implications for Robotics. As we show in Appendix A, our online hidden state representation learning method naturally extends to other platforms and environments. This suggests that our framework for situational awareness is a promising direction for adaptive methods in robotics. By learning hidden states that augment ego- and world-information, robots can handle variability and uncertainty in a wide range of tasks, thanks to reduced ambiguity. This ability to adapt to environmental and system-specific factors unlocks many possibilities without significantly adding a lot of complexity to the systems.

### v3

@@ 5.6.2 Real World Experiments:
We evaluate the zero-shot performance of our situationally-aware dynamics model by applying it directly from simulation to real-world scenarios without additional fine-tuning. Figure 15 highlights various challenging terrains successfully navigated using our method, demonstrating its robustness and adaptability in real-world conditions. The Supplementary Video 1 (SV1): https://youtu.be/7a-NzAackE4, showcases the robot successfully navigating these scenarios. Additionally in Supplementary Video 2 (SV2): https://youtu.be/Eo68c5APuAo, we also analyze the real time responsiveness of the situational awareness module as the robot traverses new scenarios.
@@
We also move to more challenging tasks with steeper and more rugged terrain, which is shown in Figure 5 (right). To protect the hardware, we only evaluated the best performing baseline in the previous task (Dubins). However, we had to stop the experiment with the Dubins model as it repeatedly led to the robot turning over and rolling down the hill, as shown in Figure 17. In contrast, our method repeatedly completed the task, as can be seen starting at the 1:42 minute mark of SV1. Overall, our method can lead the robot to successfully climb bumpy and muddy hills, whereas the baselines fail as they get stuck on terrain irregularities or are unable to devise a safe and successful plan.
@@
  • Backing Up: Whenever forward motion is impeded, we see the robot back up and try moving with a slight orientation change to overcome the terrain irregularities. This mirrors the behavior of human drivers when their vehicles become stuck in challenging terrains. Check the clips from SV1 at the: 1:00, 1:27 (trial 4), 2:06, 2:36 (trial 4), 3:30 & 6:14 (rocks) minute marks for evidence of this.
  • Velocity Control: We have seen multiple evidence of how the robot’s planned velocity commands differ according to its current situation. For example, in Figure 13 we saw, for every situation, different plans to reach the same goal from the same starting state. Moreover, at the 4:00 minute mark of the SV1, the robot maintains a slower pace while navigating downhill and then deliberately accelerates (minute 4:15) upon a situation change–reaching safer, more stable terrain.
  • Trajectory Smoothness: On very rugged and dangerous terrains the robot followed curved patterns (see the clips from SV1 at minutes 1:42, 2:22 & 3:18). On the other hand, when the terrain was not as rugged and steep, the robot approached its targets directly, relying only on the velocity control to maintain its integrity (4:00, 4:22 & 6:04 minute marks).
@@
We dive deeper into why we believe this is happening in Section 7. At a high-level, we hypothesize that upon a situation change, the optimization process carried out by the sampling-based planner will be bounded by new constraints given by the dynamics of the current situation, thus generating new diverse action plans.
Furthermore, we want to highlight that the UGV platform used (Clearpath Jackal) is a rigid-body robot with no Ackermann-style steering mechanism, minimal suspension, and relatively low ground clearance. This means that the robot often experiences contact between the vehicle body and external factors (grass, rocks, tree roots, etc.), so the chassis is not consistently clear of obstructions. These physical constraints become latent factors that can also affect the dynamics and some maneuvers that might be feasible for other vehicles can be unsafe or damaging for this platform.
@@
Figure 16: Results from the UGV experiments in the real world. We repeat the experiments ten times with each baseline and report the success rate, number of waypoints reached, the cumulative time between successive waypoints and the total time to complete the task.
@@
Figure 17: We test our situationally-aware dynamics model against the physics-based Dubins model (best performing baseline) in the challenging task of descending a steep hill with wet leaves and many large hidden bumps. Across multiple trials, our model consistently succeeded in completing the task, whereas the baseline failed, causing the robot to flip.
6 Results: Quadrotor
Figure 18: Crazyflie simulated in PyFlyt’s gates environment with changing wind conditions. The next target gate is colored green and the remaining targets yellow.
We continue to assess our situationally-aware dynamics model with a different platform type, seeking to provide further evidence to answer the three questions posited in Section 5. Specifically, we investigate the performance of SA for a quadrotor tasked with crossing a series of gates suppend in the air while being disturbed by sudden changes in wind patterns. Figure 18 presents an example of the task. Note that the agent has neither knowledge of when the change happens nor a dedicated sensor to measure specific variables that would let it determine when the wind conditions have changed or how they have changed. Therefore, it needs to rely on its proprioception to decide how to adapt.
In Section 6.1 we introduce the multiple evaluation scenarios. Section 6.2 covers the evaluation setup, including a description of the state, action and situation space. Finally, in Section 6.3 we show that leveraging a situationally-aware dynamics model the quadrotor achieves a higher reward on the task across environments with different levels of difficulty and with multiple data budgets.
6.1 Evaluation Scenarios
We leverage the PyFlyt simulator Tai et al. (2023). The main driver for our selection was the ease to implement different wind patterns, including randomly switching wind fields. In particular, we focused on the gate-crossing task, where a quadrotor needs to cross five gates suspended in the air. We modified the base task to include switching wind fields with three different levels of difficulty, but leave the reward computation and remaining components of the original task unmodified.
We designed three distinct environments, with different levels of complexity. In each environment, the active wind condition is randomly selected from a predefined set and persists for a random duration before switching to a new, different condition. The specifics of each environment are detailed in Table 4.
Environment Switching Interval (steps) Wind Type Key Parameters
1 [15, 45]
Nominal
Thermal Updraft
Strength = 1.0
2 [15, 45]
Nominal
Wind Gradient
Wind Gradient
Speed at 10m = 2.0 m/s, From 180°, $z_{0}=0.1$
Speed at 5m = 5.0 m/s, From 45°, $z_{0}=0.1$
3 [10, 50]
Nominal
Wind Gradient
Wind Gradient
Thermal Updraft
Speed at 10m = 5.0 m/s, From 270°, $z_{0}=0.3$
Speed at 3m = 5.0 m/s, From 90°, $z_{0}=0.1$
Strength = 3.0
Table 4: Wind field configurations for the quadrotor gate-crossing task. During execution, the environments randomly switch between their available wind types without revealing this information to the agent. The agent has to rely on its proprioception to detect and characterize this conditions and adapt accordingly.
6.2 Experimental Setup
6.2.1 State and Action Space:
The state space $s\in\mathbb{R}^{12}$ corresponds to $s=\begin{bmatrix}\mathbf{\Omega}&\mathbf{e}&\mathbf{v}&\mathbf{p}\end{bmatrix}^{\top}$, where $\mathbf{\Omega}=[\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{3}$ are the angular velocity components, $\mathbf{e}=[\phi,\zeta,\gamma]\in\mathbb{R}^{3}$ are the roll, pitch, and yaw angles (orientation), $\mathbf{v}=[v_{x},v_{y},v_{z}]\in\mathbb{R}^{3}$ are the linear velocity components and $\mathbf{p}=[x,y,z]\in\mathbb{R}^{3}$ is the drone’s position.
The action space is defined as $a=\begin{bmatrix}\omega^{x}_{\text{cmd}}&\omega^{y}_{\text{cmd}}&\omega^{z}_{\text{cmd}}&T\ \end{bmatrix}^{\top}\in\mathbb{R}^{4}$, where $\omega^{x}_{\text{cmd}},\omega^{y}_{\text{cmd}},\omega^{z}_{\text{cmd}}\in[-\pi,\pi]$ are the commanded angular velocities, and $T\in[0,0.8]$ is the thrust command.
6.2.2 Situation Space:
Since we defined the situation of the robot as the joint transition distribution, we concatenate the state and action variables to obtain a situation vector $\mathbf{x}\in\mathbb{R}^{28}$:
where the states $s_{t-1}$ and $s_{t}$ are preprocessed to ensure invariance and continuity as described in Section 5.3.
6.3 Situationally-Aware Dynamics Learning
To learn the quadrotor’s dynamics we leverage a dataset of state transitions collected in each of the three environments. Since no exploration by the agent needs to take place, we proceed in the passive training fashion described in Section 4.1.1. That is, the dataset is fed sequentially to the agent and the model is then trained after the situation-augmented replay buffer has been obtained.
We train two agents. One corresponds to a probabilistic ensemble (PETS) Chua et al. (2018). The other one, our agent which we call SA-PETS, is the exact same probabilistic ensemble but its input has been augmented to also include the situation symbol obtained by our online and unsupervised latent factor representation method.
Both probabilistic ensembles consist of five independently trained Gaussian multi-layer perceptrons (MLPs) with residual connections. Each network takes as input the concatenation of the current state and control vector, and outputs the mean and log-variance of a Gaussian distribution of the predicted state change. The architecture begins with a fully connected layer mapping the input dimension to a hidden size of 1024, followed by Layer Normalization and a Mish activation to stabilize training and improve gradient flow. This is followed by two residual blocks, each containing two linear layers with LayerNorm and Mish activations, where the input to the block is added back to its output and passed through an additional Mish nonlinearity. Finally, a linear projection layer maps the hidden representation to twice the dimensionality of the state space, producing both the mean and log-variance of the Gaussian distribution.
Both inputs and targets are normalized using statistics computed from the training data, and predictions are de-normalized at inference time. The model is trained using the Gaussian negative log-likelihood (NLL) loss with the Adam optimizer (learning rate $10^{-3}$). As described in Section 4.1.1, SA-PETS uses a weighted version of NLL. Training employs a learning rate scheduler (factor $0.1$, patience $10$) and early stopping with a patience of $16$ epochs. The batch size is set to 256, and the maximum number of epochs is 256. Table 5 presents the SA and MPPI hyperparameters. The sampling based controller uses the same cost function and parameters with SA-PETS and PETS.
Algorithm Hyperparameter Value
Situational Awareness
Hazard Rate $\lambda$ $60$
Threshold $\tau$ $200$
Situation Support $40$
MPPI
Number of iterations 22
Population size $1024$
Reward scaling $\gamma$ $0.9$
Noise scaling $\sigma$ $0.4$
Correlation term $\beta$ $0.4$
Table 5: Hyperparameters used for Situational Awareness and MPPI in the Quadrotor flying under changing wind task.
Figure 19: Quadrotor performance on three environments with different randomly changing wind fields. Shown are task rewards across ten episodes after training on 100k (left) and 5M (right) state transitions; higher reward indicates greater robustness.
6.3.1 Training Results:
Figure 19 presents the results after training PETS and SA-PETS with 100.000 (left) and 5 million (right) state transitions. In general, we can see SA-PETS significantly outperforms the probabilistic ensemble that does not have access to situation information.
Our proposed method, SA-PETS, demonstrates marked superiority over the PETS baseline in both data-limited and data-rich regimes, as detailed in Figure 19. In the low-data scenario with only 100k transitions, SA-PETS already achieves a substantial 34.2% overall performance improvement, showcasing significant gains in data efficiency. This advantage becomes even more pronounced as the models are exposed to more experience. With 5M transitions, the performance gap widens, culminating in a 43.6% overall improvement. Most strikingly, we observe a dramatic performance divergence in Environment 2, where the improvement leaps from 0.5% at 100k transitions to a remarkable 197.0% at 5M transitions. This results strongly indicate that adding situational awareness to the standard PETS model provides a fundamental advantage in model capacity, enabling it to capture latent factors (wind conditions in this case) and exploit complex system dynamics that the baseline model fails to identify, even when provided with large datasets. Such a characteristic is critical for achieving robust, high-performance control in complex, real-world robotic applications.
Upon closer inspection, we also notice an interesting pattern. For instance, in Environment 2, the performance of both PETS and SA-PETS actually degrades as the robot has access to more data. We hypothesize this is due to the fact that larger datasets include the presence of more latent factors —as the robot experiences more wind patterns— that confound the dynamics learning process, thereby making it harder for standard models to generalize effectively. Nonetheless, we see that the model augmented with situational awareness is significantly more robust: its task reward only decreased moderately (from 280.27 to 190.65), whereas the baseline PETS suffered a much sharper drop (from 278.83 to 64.19). This highlights not only the stabilizing effect of situational awareness in the face of noisy or confounding data but also suggests that simply collecting more data is not always beneficial—what matters is whether the model can disentangle and adapt to the hidden factors underlying the environment’s variability.
We present videos of the quadrotor flying using both dynamics models in the three environments in Supplementary Video 3 (SV3): https://youtu.be/a9oVk-Qor9c.
7 Discussion
@@
This study presented a framework for learning online and unsupervised hidden state representations to inform the dynamics models of robots operating in complex, unstructured environments. By integrating these learned representations with the dynamics models of ground and aerial vehicles we demonstrated significant performance improvements over traditional approaches that rely solely on raw sensory inputs. Particularly interesting was the emergence of safe behaviors that unlocked a new range of tasks the robot could safely accomplish. Here, we dive into the implications of these results, their connection to prior work, and the broader insights they provide for learning-based dynamics modeling in robotics.
@@
Generalization to Novel Terrains. One of the most encouraging findings was the UGV model’s ability to generalize to unseen terrains directly without additional work from simulation to reality and the emergence of safe navigation behaviors. This suggests that the learned hidden states encode features that are transferable across a broad range of terrain types. Such generalization capabilities are crucial for autonomous systems that must operate in diverse environments without extensive retraining, particularly in scenarios such as disaster response or planetary exploration.
Data Efficiency and Model Scalability. Despite the complexity of the augmented dynamics model, it demonstrated significant data efficiency compared to baseline methods. By leveraging the representations learned online, the UGV’s model required fewer training episodes to achieve superior performance, as evidenced in Figures 7 & 10. Similar results were observed in the quadrotor’s case, as demonstrated in Figure 19. This efficiency highlights the potential for scaling the approach to larger, more complex systems, and even online dynamics learning tasks, since we abandon the need for privileged information and extensive offline training from previous works on auxiliary representation learning.
@@
On the Situation Change Responsiveness. Quantitatively measuring detection lag for an unobserved latent context is inherently difficult, since ground-truth labels for latent factor changes are not available. To address responsiveness within these limits, we rely on complementary qualitative and diagnostic analyses: representative changepoints detected during operation (Figure 8 and Appendix B.1) and a 2D projection of the learned 26-dimensional local-transition descriptors that reveals distinct situation clusters (Figure 9). We also recorded three new videos in simulation and the real world (SV2: https://youtu.be/Eo68c5APuAo) where the accompanying dashboard let us visualize the system responsiveness to a situation change as well as the situation histories. There we can also appreciate the emergence of behavior such as the situation-dependent velocity control when entering more difficult terrain (as also show in Figure 13). Together, these materials show that the module segments local transition dynamics and that the learned representations encode different transition patterns rather than collapsing to a single generic embedding.
How quickly the system switches situations depends on deliberate design trade-offs in the BOCD’s changepoint logic. Two elements are particularly important: the changepoint prior, which controls how readily the model hypothesizes a switch (a higher prior yields faster detections at the cost of more false positives), and the minimum-support requirement used before instantiating a new situation (included to prevent numerical issues during precision-matrix inversion), which introduces a short, controlled detection delay. Whether a single state–action sample triggers a switch therefore depends on how out-of-distribution that sample is: in BOCD the run-length posterior depends on the likelihood of the latest observation (see Eqs. 4 and 7), so a very low likelihood can rapidly shift posterior mass toward a changepoint hypothesis. In practice, with our chosen hazard and minimum-support settings the system favors robust detection of sustained changes rather than flipping on isolated, noisy samples.
@@
Integration with Planning and Control. The improved dynamics predictions directly impacted downstream planning and control tasks, leading to safer and adaptive navigation/flying. As shown in Figures 15, 16, 17 & 19, the performance improvement is notably higher in challenging scenarios regardless of their real or simulated nature. By augmenting the state space with the hidden state representations, the controller could generate trajectory plans that proactively accounted for terrain- and wind-induced dynamics changes.
Comparison with Learning-Based Models. The performance of the CaDM and PE+LSTM baselines highlights the inherent difficulty of learning hidden state representations alongside dynamics in a data-efficient manner. While these methods have shown promise in other scenarios, our online hidden state representation framework, grounded in multivariate BOCD, offers a more effective alternative. By leveraging structured latent spaces and online learning, our approach eliminates the need for interdependent, separately trained models, addressing the problem from a novel perspective. We attribute these performance gains to the online nature of our method, which mitigates the non-stationarity in context representations that arises when dynamics and context encoders are trained simultaneously, as in the case of PE+LSTM. Moreover, by avoiding the need to learn a separate context model, our approach reduces data requirements and maintains a smaller overall model size. As discussed in Section 2.3, other methods that leverage auxiliary representations usually have to rely on multi-stage approaches and privileged environment information in order to train successful policies.
@@
On the dynamics learning side, the PE-Y model, despite using and predicting the same variables as the Dubins model, exhibited poor real-world performance. Similarly, PE-RPY struggled to generalize to new terrains; its requirement to predict the robot’s orientation (roll, pitch, and yaw) introduced compounding errors during model-based planning, which led to suboptimal strategies, as demonstrated in minute 0:51 (PE-RPY) of the SV1. We further hypothesize that TD-MPC2’s slower learning is attributable to its large model size, necessitating the simultaneous learning of five components (encoder, latent dynamics, reward predictor, terminal value predictor and policy), which may introduce unnecessary overhead for this specific task. Finally, while the model-free alternatives predictably underperformed in terms of learning efficiency, they underscore the advantages of a model-based approach for adaptive robot behavior. This is particularly compelling when contrasted with the policy-learning strategies for adaptive robot behaviors employed by other works, as discussed in Section 2.3.
@@
Comparison with Physics-Based Models. The Dubins model was the most effective baseline but with critical flaws in complex real world scenarios. Note that its performance stems from being designed by an expert with extensive domain knowledge. This adds some limitations due to the stringent prerequisites to replicate this approach for other robots. Additionally, it operates under the assumption of ideal conditions, which limits its applicability in real-world, unstructured terrain navigation tasks. Critically, the model lacks a feedback mechanism to account for the specific characteristics of varying terrains, making it ill-suited for dynamic and complex environments. This limitation was evident in several scenarios where the robot’s performance was unsatisfactory. For instance, the model was unable to compensate for latent factors affecting the robot’s motion—as captured in minute 0:37 (Dubins) of SV1—and frequently resulted in unsafe behaviors, both for the robot and its surroundings, as demonstrated in Figure 17.
On the Role of Exteroceptive Perception. Integrating exteroceptive perception (e.g., camera or LiDAR) into our situational-awareness module offers a promising route to both accelerate and make more proactive the adaptation process. Rather than replacing the latent modeling core, perceptual cues can be used to form an informed prior over changepoint probabilities—replacing the constant hazard function used in Section 4.2.2—and thereby adaptively bias the model to expect particular transitions (e.g., entering gravel or grass) before they fully manifest in dynamics. This perceptually-informed prior reduces the representational burden on the SA module (it no longer needs to infer gross, easily observable factors from dynamics alone) and enables the planner to adopt anticipatory strategies instead of purely reactive ones. Perception can also be coupled via existing traversability planners to provide intermediate objectives or risk estimates that guide the model-based controller. Important caveats for future work include robustly handling noisy or out-of-distribution sensor inputs and explicitly representing perceptual uncertainty so that incorrect priors do not degrade performance; principled fusion methods (e.g., uncertainty-aware Bayesian updates or learned perceptual priors) are natural directions. Overall, exteroceptive perception acts as a complementary information channel that can make latent adaptation faster and planning more proactive while leaving the SA module responsible for modeling residual, unmeasured latent factors (e.g., wet vs. dry grass).
@@
Broader Implications for Robotics. As we show in Section 5 (ground), Section 6 (aerial) & Appendix B (aquatic), our online hidden state representation learning method naturally extends to other platforms and environments. This suggests that our framework for situational awareness is a promising direction for adaptive methods, specially in field robotics applications. By learning hidden states that augment ego- and world-information, robots can handle variability and uncertainty in a wide range of tasks, thanks to reduced ambiguity. This ability to adapt to environmental and system-specific factors unlocks many possibilities without significantly adding a lot of complexity to the systems.
7.1 Limitations and Scope